# GVE-Leiden: Fast Leiden Algorithm for Community Detection in Shared Memory Setting

### v6

@@ 4.2.2. Local-moving phase of GVE-Leiden
Lines 12-24 represent the main loop of the local-moving phase. In line 11, we first mark all vertices as unprocessed. Then, in line 13, we initialize the total delta-modularity per iteration $\Delta Q$. Next, in lines 14-23, we iterate over unprocessed vertices in parallel. For each unprocessed vertex $i$, we mark $i$ as processed - vertex pruning (line 15), scan communities connected to $i$ excluding self (line 16), determine the best community $c*$ to move $i$ to (line 18), and calculate the delta-modularity of moving $i$ to $c*$ (line 19). We then update the community membership of $i$ (lines 21-22) and mark its neighbors as unprocessed (line 23) if a better community was found. In line 24, we check if the local-moving phase has converged. If so, we break out of the loop (or if $MAX\_ITERATIONS$ is reached). At the end, in line 25, we return the number of iterations performed $l_{i}$.
@@ 4.2.3. Refinement phase of GVE-Leiden
Lines 10-20 represent the core of the refinement phase. In the refinement phase, we perform, what is called the constrained merge procedure (Traag et al., 2019). The idea here is to allow vertices, within each community bound, to form sub-communities by only allowing isolated vertices (i.e., vertices belonging to their own community) to change their community membership. This procedure splits any internally-disconnected communities identified during the local-moving phase, and prevents the formation of any new disconnected communities. Here, for each isolated vertex $i$ (line 12), we scan communities connected to $i$ within the same community bound - excluding self (line 13), evaluate the best community $c*$ to move $i$ to (line 15), and compute the delta-modularity of moving $i$ to $c*$ (line 16). If a better community was found, we attempt to update the community membership of $i$ if it is still isolated (lines 18-20).
@@ 4.2.4. Aggregation phase of GVE-Leiden
Then, in lines 18-24, a parallel iteration over all communities $c\in[0,|\Gamma|)$ is performed. For each vertex $i$ belonging to community $c$, all communities $d$ (with associated edge weight $w$), linked to $i$ as defined by scanCommunities() in Algorithm 2, are added to the per-thread hashtable $H_{t}$. Once $H_{t}$ is populated with all communities (and associated weights) linked to community $c$, these are atomically added as edges to super-vertex $c$ in the super-vertex graph $G^{\prime\prime}$. Finally, in line 25, we return the super-vertex graph $G^{\prime\prime}$.
@@ 5.1.1. System used
We employ a server equipped with two Intel Xeon Gold 6226R processors, each featuring $16$ cores running at a clock speed of $2.90$ GHz. Each core is equipped with a $1$ MB L1 cache, a $16$ MB L2 cache, and a $22$ MB shared L3 cache. The system is configured with $376$ GB RAM and set up with CentOS Stream 8.
@@ 5.1.2. Configuration
We use 32-bit integers for vertex ids and 32-bit float for edge weights but use 64-bit floats for computations and hashtable values. We utilize $64$ threads to match the number of cores available on the system (unless specified otherwise). For compilation, we use GCC 8.5 and OpenMP 4.5.
@@ 5.1.3. Dataset
(a) Runtime in seconds (logarithmic scale) with Original Leiden, igraph Leiden, NetworKit Leiden, and GVE-Leiden
(b) Speedup of GVE-Leiden (logarithmic scale) with respect to Original Leiden, igraph Leiden, NetworKit Leiden.
(c) Modularity of communities obtained with Original Leiden, igraph Leiden, NetworKit Leiden, and GVE-Leiden.
(d) Fraction of disconnected communities (logarithmic scale) with Original Leiden, igraph Leiden, NetworKit Leiden, and GVE-Leiden.
Figure 4. Runtime in seconds (log-scale), speedup (log-scale), modularity, and fraction of disconnected communities (log-scale) with Original Leiden, igraph Leiden, NetworKit Leiden, and GVE-Leiden for each graph in the dataset.
@@ 5.2. Comparing Performance of GVE-Leiden
We now compare the performance of GVE-Leiden with the original Leiden (Traag et al., 2019), igraph Leiden (Csardi et al., 2006), and NetworKit Leiden (Staudt et al., 2016). For the original Leiden, we use a C++ program to initialize a ModularityVertexPartition upon the loaded graph, and invoke optimise_partition() to obtain the community membership of each vertex in the graph. On graphs with a large number of edges, such as webbase-2001 and sk-2005, using ModularityVertexPartit ion introduces disconnected communities due to issues with numerical precision (i.e. the improvement of separating two disconnected parts may be positive, but due to the enormous weight, this may effectively be near 0) (Traag, 2024). For such graphs, we instead use RBConfigurationVertexPartition, which uses unscaled improvements to modularity (i.e. they do not scale with the total weight). For igraph Leiden, we use igraph_community_leiden() with a resolution of $1/2|E|$, a beta of $0.01$, and request the algorithm to run until convergence. For NetworKit Leiden, we write a Python script to call ParallelLeiden(), while limiting the number of passes to $10$. For each graph, we measure the runtime of each implementation and the modularity of the communities obtained, five times, for averaging. We also save the community membership vector to a file and later count the number of disconnected components using Algorithm 5. In all instances, we use modularity as the quality function to optimize for.
@@
Figure 4(a) shows the runtimes of the original Leiden, igraph Leiden, NetworKit Leiden, and GVE-Leiden on each graph in the dataset. On the sk-2005 graph, GVE-Leiden finds communities in $9.4$ seconds, and thus achieve a processing rate of $403$ million edges/s. Figure 4(b) shows the speedup of GVE-Leiden with respect to each implementation mentioned above. GVE-Leiden is on average 436×436$\times$436 ×, 104×104$\times$104 ×, and 8.2×8.2$\times$8.2 × faster than the original Leiden, igraph Leiden, and NetworKit Leiden respectively. Figure 4(c) shows the modularity of communities obtained with each implementation. GVE-Leiden on average obtains $0.3\%$ lower modularity than the original Leiden and igraph Leiden, and $25\%$ higher modularity than NetworKit Leiden (especially on road networks and protein k-mer graphs). Finally, Figure 4(d) shows the fraction of disconnected communities obtained with each implementation. Here, the absence of bars indicates the absence of disconnected communities. Communities identified by NetworKit Leiden have on average $1.5\times 10^{-2}$ fraction of disconnected communities, while none of the communities identified by the original Leiden, igraph Leiden, and GVE-Leiden are internally-disconnected. As the Leiden algorithm guarantees the absence of disconnected communities (Traag et al., 2019), those observed with NetworKit Leiden are likely due to implementation issues.
Next, we compare the performance of GVE-Leiden with GVE-Louvain (Sahu, 2023), our parallel implementation of the Louvain method. As above, for each graph in the dataset, we run both algorithms 5 times to minimize measurement noise. Figure 5(a) shows the runtimes of GVE-Louvain and GVE-Leiden on each graph in the dataset. Figure 5(b) shows the speedup of GVE-Leiden with respect to GVE-Louvain. GVE-Leiden is on average $13\%$ slower than GVE-Louvain. This increase in computation time is a trade-off for identifying communities that are not internally-disconnected, as given below. Figure 5(c) shows the modularity of communities obtained with GVE-Louvain and GVE-Leiden. GVE-Leiden on average obtains the same modularity as GVE-Louvain. Finally, Figure 5(d) shows the fraction of internally-disconnected communities obtained. Communities identified by GVE-Louvain on average have $4.0\%$ disconnected communities, while GVE-Leiden has none.
@@
Figure 7. Runtime $/|E|$ factor with GVE-Leiden for each graph in the dataset.
@@ 5.3. Analyzing Performance of GVE-Leiden
We now analyze the performance of GVE-Leiden. The phase-wise and pass-wise split of GVE-Leiden is shown in Figures 6(a) and 6(b) respectively. Figure 6(a) reveals that GVE-Leiden devotes a significant portion of its runtime to the local-moving and refinement phases on web graphs, road networks, and protein k-mer graphs, while it dedicates majority of its runtime in the aggregation phase on social networks. The pass-wise split, shown in Figure 6(b), indicates that the first pass is time-intensive for high-degree graphs (web graphs and social networks), while subsequent passes take precedence in execution time on low-degree graphs (road networks and protein k-mer graphs).
@@ 6. Conclusion
In conclusion, this study addresses the design of the most optimized multicore implementation of the Leiden algorithm (Traag et al., 2019), to the best of our knowledge. On a system equipped with two 16-core Intel Xeon Gold 6226R processors, our implementation of the Leiden algorithm, referred to as GVE-Leiden, attains a processing rate of $403M$ edges per second on a $3.8B$ edge graph. It surpasses the original Leiden implementation, igraph Leiden, and NetworKit Leiden by factors of 436×436$\times$436 ×, 104×104$\times$104 ×, and 8.2×8.2$\times$8.2 × respectively. GVE-Leiden identifies communities of equivalent quality to the first two implementations, and $25\%$ higher quality than NetworKit. Doubling the number of threads results in an average performance scaling of 1.6×1.6$\times$1.6 × for GVE-Leiden. In comparison to GVE-Louvain (our parallel Louvain implementation) (Sahu, 2023), the original Leiden, igraph Leiden, and NetworKit Leiden, GVE-Leiden completely eliminates internally-disconnected communities.
@@ Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Vincent Traag, Geerten Verweij, and Fabian Nguyen for their support.
@@ Appendix A Appendix
A.1. Community labels of super-vertices
We also attempt two different variations of Parallel Leiden algorithm, one where the community labels of super-vertices (upon aggregation) is based on the local-moving phase (move-based), and the other where the community labels of super-vertices is based on the refinement phase (refine-based). Our observations indicate that both approaches have roughly the same runtime and modularity on average, as indicated by Figures 9 and 10. Accordingly, we stick to the move-based approach, which is the one recommended by Traag et al. (Traag et al., 2019). However, refine-based approach may be more suitable for the design of dynamic Leiden algorithm (for dynamic graphs).
A.2. Finding disconnected communities
@@
A.3. Indirect Comparison with State-of-the-art Leiden Implementations
Finally, we conduct an indirect comparison of the performance of our multicore implementation of the Leiden algorithm with other similar state-of-the-art implementations, as listed in Table 1. Hu et al. (Hu et al., [n. d.]) introduce ParLeiden, a parallel Leiden implementation for distributed environments, which uses thread locks and efficient buffers, to resolve community joining conflicts and reduce communication overheads. They refer to their single node version of ParLeiden as ParLeiden-S, and their distributed version as ParLeiden-D. On a cluster with $8$ nodes, with each node being equipped with a $48$ core CPU, Hu et al. observe a speedup of 12.3×12.3$\times$12.3 ×, $\times9$, and 1.32×1.32$\times$1.32 × for ParLeiden-S, ParLeiden-D, and a baseline Leiden implemented on KatanaGraph, on the com-LiveJournal graph, with respect to original Leiden (Traag et al., 2019) (refer to Table 2 in their paper (Hu et al., [n. d.])). In contrast, on the same graph, we observe a speedup of 219×219$\times$219 × relative to original Leiden. Consequently, our Leiden implementation outperforms ParLeiden-S, ParLeiden-D, and KatanaGraph Leiden by approximately 18×18$\times$18 ×, $\times2$, and 166×166$\times$166 × respectively.

### v7

@@ 4.2.2. Local-moving phase of GVE-Leiden
Lines 11-23 represent the main loop of the local-moving phase. In line 10, we first mark all vertices as unprocessed. Then, in line 12, we initialize the total delta-modularity per iteration $\Delta Q$. Next, in lines 13-22, we iterate over unprocessed vertices in parallel. For each unprocessed vertex $i$, we mark $i$ as processed - vertex pruning (line 14), scan communities connected to $i$ excluding self (line 15), determine the best community $c*$ to move $i$ to (line 17), and calculate the delta-modularity of moving $i$ to $c*$ (line 18). We then update the community membership of $i$ (lines 20-21) and mark its neighbors as unprocessed (line 22) if a better community was found. In line 23, we check if the local-moving phase has converged. If so, we break out of the loop (or if $MAX\_ITERATIONS$ is reached). At the end, in line 24, we return the number of iterations performed $l_{i}$.
@@ 4.2.3. Refinement phase of GVE-Leiden
Lines 8-17 represent the core of the refinement phase. In the refinement phase, we perform, what is called the constrained merge procedure (Traag et al., 2019). The idea here is to allow vertices, within each community bound, to form sub-communities by only allowing isolated vertices (i.e., vertices belonging to their own community) to change their community membership. This procedure splits any internally-disconnected communities identified during the local-moving phase, and prevents the formation of any new disconnected communities. Here, for each isolated vertex $i$ (line 10), we scan communities connected to $i$ within the same community bound - excluding self (line 11), evaluate the best community $c*$ to move $i$ to (line 13), and compute the delta-modularity of moving $i$ to $c*$ (line 14). If a better community was found, we attempt to update the community membership of $i$ if it is still isolated (lines 16-17).
@@ 4.2.4. Aggregation phase of GVE-Leiden
Then, in lines 18-23, a parallel iteration over all communities $c\in[0,|\Gamma|)$ is performed. For each vertex $i$ belonging to community $c$, all communities $d$ (with associated edge weight $w$), linked to $i$ as defined by scanCommunities() in Algorithm 2, are added to the per-thread hashtable $H_{t}$. Once $H_{t}$ is populated with all communities (and associated weights) linked to community $c$, these are atomically added as edges to super-vertex $c$ in the super-vertex graph $G^{\prime\prime}$. Finally, in line 24, we return the super-vertex graph $G^{\prime\prime}$.
@@ 5.1.1. System used
We employ a server equipped with two Intel Xeon Gold 6226R processors, each featuring $16$ cores running at a clock speed of $2.90$ GHz. Each core is equipped with a $1$ MB L1 cache, a $16$ MB L2 cache, and a $22$ MB shared L3 cache. The system is configured with $376$ GB RAM and set up with CentOS Stream 8. For our GPU experiments, we utilize a system with an NVIDIA A100 GPU (108 SMs, 64 CUDA cores per SM, 80 GB global memory, 1935 GB/s bandwidth, 164 KB shared memory per SM) and an AMD EPYC-7742 processor (64 cores, 2.25 GHz). The server is equipped with 512 GB DDR4 RAM and runs Ubuntu 20.04.
@@ 5.1.2. Configuration
We use 32-bit integers for vertex ids and 32-bit float for edge weights but use 64-bit floats for computations and hashtable values. We utilize $64$ threads to match the number of cores available on the system (unless specified otherwise). For compilation, we use GCC 8.5 and OpenMP 4.5 on the CPU system, and GCC 9.4, OpenMP 5.0, and CUDA 11.4 on the GPU system.
@@ 5.1.3. Dataset
(a) Runtime in seconds (logarithmic scale) with Original Leiden, igraph Leiden, NetworKit Leiden, cuGraph Leiden, and GVE-Leiden
(b) Speedup of GVE-Leiden (logarithmic scale) with respect to Original Leiden, igraph Leiden, NetworKit Leiden, and cuGraph Leiden.
(c) Modularity of communities obtained with Original Leiden, igraph Leiden, NetworKit Leiden, cuGraph Leiden, and GVE-Leiden.
(d) Fraction of disconnected communities (logarithmic scale) with Original Leiden, igraph Leiden, NetworKit Leiden, cuGraph Leiden, and GVE-Leiden.
Figure 6. Runtime in seconds (log-scale), speedup (log-scale), modularity, and fraction of disconnected communities (log-scale) with Original Leiden, igraph Leiden, NetworKit Leiden, cuGraph Leiden, and GVE-Leiden for each graph in the dataset.
@@ 5.2. Comparing Performance of GVE-Leiden
We now compare the performance of GVE-Leiden with the original Leiden (Traag et al., 2019), igraph Leiden (Csardi et al., 2006), NetworKit Leiden (Staudt et al., 2016), and cuGraph Leiden (Kang et al., 2023). The original Leiden and igraph Leiden are sequential implementations, while NetworKit Leiden and GVE-Leiden are parallel multicore implementations of the Leiden algorithm. On the other hand, cuGraph Leiden is a GPU implementation of Leiden, and is run on our GPU-based system. For the original Leiden, we use a C++ program to initialize a ModularityVertexPartition upon the loaded graph, and invoke optimise_partition() to obtain the community membership of each vertex in the graph. On graphs with a large number of edges, such as webbase-2001 and sk-2005, using ModularityVertexPartition introduces disconnected communities due to issues with numerical precision (i.e. the improvement of separating two disconnected parts may be positive, but due to the enormous weight, this may effectively be near 0) (Traag, 2024). For such graphs, we instead use RBConfigurationVertexPartition, which uses unscaled improvements to modularity (i.e. they do not scale with the total weight). For igraph Leiden, we use igraph_comm unity_leiden() with a resolution of $1/2|E|$, a beta of $0.01$, and request the algorithm to run until convergence. For NetworKit Leiden, we write a Python script to call ParallelLeiden(), while limiting the number of passes to $10$. For cuGraph Leiden, we write a Python script to configure the Rapids Memory Manager (RMM) to use a pool allocator for fast memory allocations and run cugraph.leiden() on the loaded graph. For each graph, we measure the runtime of each implementation and the modularity of the communities obtained, five times, for averaging. With cuGraph, we disregard the runtime of the first run to ensure subsequent measurements only reflect RMM’s pool usage without CUDA memory allocation overhead. We save the obtained community membership vector to a file and later count the disconnected communities using an algorithm given in our extended report (Sahu, 2023a). In all instances, we use modularity as the quality function to optimize for.
@@
Figure 6(a) shows the runtimes of the original Leiden, igraph Leiden, NetworKit Leiden, cuGraph Leiden, and GVE-Leiden on each graph in the dataset. cuGraph Leiden fails to run on the arabic-2005, uk-2005, webbase-2001, it-2004, and sk-2005 graphs due to out of memory issues. On the sk-2005 graph, GVE-Leiden finds communities in $9.4$ seconds, and thus achieve a processing rate of $403$ million edges/s. Figure 6(b) shows the speedup of GVE-Leiden with respect to each implementation mentioned above. GVE-Leiden is on average 436×436$\times$436 ×, 104×104$\times$104 ×, 8.2×8.2$\times$8.2 ×, and 3.0×3.0$\times$3.0 × faster than the original Leiden, igraph Leiden, NetworKit Leiden, and cuGraph Leiden respectively. Figure 6(c) shows the modularity of communities obtained with each implementation. GVE-Leiden on average obtains $0.3\%$ lower modularity than the original Leiden and igraph Leiden, $25\%$ higher modularity than NetworKit Leiden (especially on road networks and protein k-mer graphs), and $3.5\%$ higher modularity that cuGraph Leiden (primarily due to cuGraph Leiden’s inability to run on well-clusterable graphs). Finally, Figure 6(d) shows the fraction of disconnected communities obtained with each implementation. Here, the absence of bars indicates the absence of disconnected communities. Communities identified by NetworKit Leiden and cuGraph Leiden have on average $1.5\times 10^{-2}$ and $6.6\times 10^{-5}$ fraction of disconnected communities, respectively, while none of the communities identified by the original Leiden, igraph Leiden, and GVE-Leiden are internally-disconnected. As the Leiden algorithm guarantees the absence of disconnected communities (Traag et al., 2019), those observed with NetworKit Leiden and cuGraph Leiden are likely due to implementation issues.
@@
Figure 9. Runtime $/|E|$ factor with GVE-Leiden for each graph.
@@
Next, we compare the performance of GVE-Leiden with GVE-Louvain (Sahu, 2023b), our parallel implementation of the Louvain method. As above, for each graph in the dataset, we run both algorithms 5 times to minimize measurement noise. Figure 7(a) shows the runtimes of GVE-Louvain and GVE-Leiden on each graph in the dataset. Figure 7(b) shows the speedup of GVE-Leiden with respect to GVE-Louvain. GVE-Leiden is on average $13\%$ slower than GVE-Louvain. This increase in computation time is a trade-off for identifying communities that are not internally-disconnected, as given below. Figure 7(c) shows the modularity of communities obtained with GVE-Louvain and GVE-Leiden. GVE-Leiden on average obtains the same modularity as GVE-Louvain. Finally, Figure 7(d) shows the fraction of internally-disconnected communities obtained. Communities identified by GVE-Louvain on average have $4.0\%$ disconnected communities, while GVE-Leiden has none.
@@ 5.3. Analyzing Performance of GVE-Leiden
We now analyze the performance of GVE-Leiden. The phase-wise and pass-wise split of GVE-Leiden is shown in Figures 8(a) and 8(b) respectively. Figure 8(a) reveals that GVE-Leiden devotes a significant portion of its runtime to the local-moving and refinement phases on web graphs, road networks, and protein k-mer graphs, while it dedicates majority of its runtime in the aggregation phase on social networks. The pass-wise split, shown in Figure 8(b), indicates that the first pass is time-intensive for high-degree graphs (web graphs and social networks), while subsequent passes take precedence in execution time on low-degree graphs (road networks and protein k-mer graphs).
@@ 6. Conclusion
In conclusion, this study addresses the design of the most optimized multicore implementation of the Leiden algorithm (Traag et al., 2019), to the best of our knowledge. Here, we extend optimizations from our implementation of the Louvain algorithm (Sahu, 2023b), and use a greedy refinement phase where vertices greedily optimize for delta-modularity within their community bounds, which we observe, offers both better performance and quality than a randomized approach. On a system equipped with two 16-core Intel Xeon Gold 6226R processors, our implementation of the Leiden algorithm, referred to as GVE-Leiden, attains a processing rate of $403M$ edges per second on a $3.8B$ edge graph. It outperforms the original Leiden implementation, igraph Leiden, NetworKit Leiden, and cuGraph Leiden (run on an NVIDIA A100 GPU) by 436×436$\times$436 ×, 104×104$\times$104 ×, 8.2×8.2$\times$8.2 ×, and 3.0×3.0$\times$3.0 × respectively. GVE-Leiden identifies communities of equivalent quality to the first two implementations, and $25\%$ / $3.5\%$ higher quality than NetworKit / cuGraph. Doubling the number of threads results in a average performance scaling of 1.6×1.6$\times$1.6 × for GVE-Leiden.
@@ Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Vincent Traag, Geerten Verweij, Fabian Nguyen, Chuck Hastings, and Rick Ratzel for their support.
@@ Appendix A Appendix
A.1. Finding disconnected communities
@@
A.2. Indirect Comparison with State-of-the-art Leiden Implementations
Finally, we conduct an indirect comparison of the performance of our multicore implementation of the Leiden algorithm with other similar state-of-the-art implementations, as listed in Table 1. Please consider the reported speedups as approximate. Hu et al. (Hu et al., [n. d.]) introduce ParLeiden, a parallel Leiden implementation for distributed environments, which uses thread locks and efficient buffers, to resolve community joining conflicts and reduce communication overheads. They refer to their single node version of ParLeiden as ParLeiden-S, and their distributed version as ParLeiden-D. On a cluster with $8$ nodes, with each node being equipped with a $48$ core CPU, Hu et al. observe a speedup of 12.3×12.3$\times$12.3 ×, $\times9$, and 1.32×1.32$\times$1.32 × for ParLeiden-S, ParLeiden-D, and a baseline Leiden implemented on KatanaGraph, on the com-LiveJournal graph, with respect to original Leiden (Traag et al., 2019) (refer to Table 2 in their paper (Hu et al., [n. d.])). In contrast, on the same graph, we observe a speedup of 219×219$\times$219 × relative to original Leiden. Consequently, GVE-Leiden outperforms ParLeiden-S, ParLeiden-D, and KatanaGraph Leiden by approximately 18×18$\times$18 ×, $\times2$, and 166×166$\times$166 × respectively.